One year, the price of a flat was 127365 €, and the consumer price index (CPI) was 105.8. A couple of years later, the price has risen to 135149 €, and the consumer price index (CPI) was 112.2. Calculate the relative real change in the price as percentage.


1
Expert's answer
2022-02-10T16:07:59-0500

Rate of inflation is: "(112.2 - 105.8)\/105.8 = 0.0605" or 6.05%.

The relative nominal change in her salary is:

"(135,149 - 127,365)\/127,365 = 0.0611" or 6.11%.

The relative real change in her salary (purchasing power) as percentage is:

6.11 - 6.05 = 0.06%.